# A Spatiotemporal, Quasi-experimental Causal Inference Approach to Characterize the Effects of Global Plastic Waste Export and Burning on Air Quality Using Remotely Sensed Data

### v3

@@ 1 Introduction
Our proposed method merges ideas from this literature on quasi-experiments with a “universal” intervention and ideas from the burgeoning literature on multiply-robust causal exposure-response curve estimation. Specifically, we build closely upon recent work by Hettinger et al. (2025a), who introduced a multiply-robust estimator for causal exposure-response curves within the difference-in-differences (DiD) framework. Derived from the efficient influence function or EIF (see Section 3.2), this estimator accounts for confounding both between treated and control units and across levels of exposure to the treatment / intervention, which can be thought of as a dosage. We propose a related EIF-based multiply-robust estimator for causal exposure-response curves in quasi-experimental designs with a universal intervention. We also propose an uncertainty quantification approach that accounts for both spatial and temporal correlation. We apply this method to remotely sensed data for the Indonesia plastic waste case study to estimate the change in air pollution at open dump sites due to increased plastic waste in the wake of China’s 2018 ban, as a function of a dump site’s port proximity.
@@ 2.1.1 Why Non-dump Sites Should Not Be Used as “Control” Locations
To motivate our methodological approach (in which we use past years as controls for future years at the GPW dump sites), briefly consider what trying to identify separate “control” locations would look like. Let us walk through several of the causal identifying assumptions made in Hettinger et al. (2025a), which are standard in DiD-type analyses.
@@
We extract these monthly PM2.5 values for 2012-2019 at the GPW sites, for each site taking the value of the grid cell in which the site centroid lies. The average PM2.5 concentration at the dump sites in 2012-2017 (prior to the China ban) was 25.7 $\mu g/m^{3}$.
@@ 3.1 Setup and Assumptions
Let $\mathcal{O}$ be the observed data, containing: $Y$ the outcome (PM2.5), $D$ the dose of exposure (port proximity), and $\mathbf{X}$ the covariates. The latter includes the population living within 1km of the site, monthly averaged meteorology (temperature, humidity, precipitation, pressure, and wind speed), and the number of active fires detected in the province (that month) which did not overlap with any GPW site. To help capture spatial heterogeneity, we also include fixed effects for site ID and province in our model; the spatial boundaries for the latter are shown in Figure S3. This is considered a reasonable approach in settings with many units where only few of these are important/informative (Fuhr and Papies, 2024) and when both bigger-scale (e.g., province level) and smaller-scale (individual dump site level) heterogeneity are likely at play. Both of these characteristics are present in our case study, as shown by a complementary exploratory analysis (described in Section 4.2) illustrating that the distribution of the number of large fires (detectable by VIIRS) occurring at each dump site is highly skewed, with some clustering by province. Lastly, to account for longer-term trends, we include time t=1,,96t=1,...,96 to represent the months between 2012 and 2019, inclusive.
@@
In order to avoid structural ill-definition, we also introduce an indicator for $t\geq T_{0}$, where $T_{0}$ is the treatment initiation time. In practice, $T_{0}$ is 1/1/2018 for all sites, but conceptually this could have occurred at any point in our study period, such that any time point could have been untreated.
Our aim is to estimate a causal exposure-response curve quantifying the causal effect on Indonesian air quality of China’s 2018 plastic waste import ban at different levels of port proximity, $\delta$. In the DiD setting, Hettinger et al. (2025a) terms this the “Average Dose Effect on the Treated”, or ADT. While noting again that our formulation is different, we still see fit to call our effect of interest the ADT:
@@
An important note about the ADT is how it differs from a more commonly-used Conditional Average Treatment Effect, $CATE(\delta):=E[Y_{it}(B=1)-Y_{it}(B=0)|D_{it}=\delta]$. First, recall that $D$ is undefined when $B=0$, not yet having been ‘activated’. Second, foreshadowing the following sections, the ADT approach enables interpreting the final effect curve in terms of port proximity versus in terms of port proximity-possibly-confounded-by-other-factors. It does so by estimating a generalized propensity score for $D$ and a binary propensity score for $B$, whereas a traditional doubly-robust estimator for the CATE would only use the latter.
@@
$\>\>\>\>\text{(i) }\epsilon\leq\pi_{B}(\mathbf{x})<1\>\forall\>\mathbf{x}\in\mathbb{X},\>\>\>\>\>\>\>\>\pi_{B}(\mathbf{x})=P(B=1|\mathbf{x})$
$\>\>\>\>\text{(ii) }\epsilon\leq\pi_{D}(\delta|\mathbf{x},B=1)\>\forall\>\mathbf{x}\in\mathbb{X}|B=1,\delta\in\mathbb{D},\>\>\>\>\>\>\>\>\pi_{D}(\delta|\mathbf{x},B=1)=p(D=\delta|B=1,\mathbf{X}=\mathbf{x})$
@@
This means that there is no unmeasured time-varying confounding – no unmeasured factor causing air pollution to be systematically different in the periods before and after China’s ban, after accounting for confounding factors. Assembling an adequate set of covariates to justify this assumption relies on domain knowledge. Its plausibility can be strengthened by adjusting for seasonal and/or long-term trends, e.g., through inclusion of a time variable in the adjustment set. Note that in this setting, we do not have to worry about time-varying confounding affected by past exposure, because China’s ban being enacted should not affect confounders like meteorology and population density in Indonesia. However, we must pay careful attention to any other policies with the potential to make air pollution, conditional on covariates, systematically different in the $B=0$ and $B=1$ time periods – see discussion of this in Section 5.
@@
Conditional ignorability (in both A5 and A6) is a stronger assumption than parallel trends, in a setting where control locations are available. However, for our application, the assumption that some locations represent the control condition post-treatment (China ban) would also be strong and likely inaccurate, as described in Section 2.1.1. Note that while we believe it is reasonable to assume conditional ignorability on distributions here given our domain knowledge, all that is needed for identifiability of our estimand is (a weaker) conditional ignorability in expectations, as is shown below.
@@ 3.2 Motivations for Efficient Influence Function-Based Estimation
where $f_{\phi}(D)$ is a flexible (nonparametric) function, with parameters $\phi$. Note that a simplified version of this where $f_{\phi}(D)=\phi D$ has been extensively analyzed, for instance see Callaway et al. (2024). There are several drawbacks of this approach. To begin, even if we allow a flexible relationship between $D$ and $Y$, we are not allowing for flexible relationships between $\mathbf{X}$ and $Y$, nor interactions between the components of $\mathbf{X}$. Importantly, we are not addressing confounding between the dose $D$ and trends in the outcome – for instance, if the overall (or seasonal) trend in air pollution is affected by population density, which is static in our dataset and which we know to be correlated with $D$ (see Section 4.2). To address this, we could make a much more complicated linear regression (e.g., by using splines and more interaction terms), or use a nonparametric machine learning algorithm to model $Y$ given $D,\mathbf{X},B$. However, this outcome model might still be misspecified. This motivates the search for a doubly-robust estimator, which incorporates both a model for the outcome and a model for the propensity score (the probability of receiving treatment), which explicitly addresses confounding, and will produce results that are consistent (reliable in a probabilistic sense) if at least one of these models is correctly specified.
We turn to estimators based on the the Efficient Influence Function (EIF). Without going into the technical details (interested readers should see Kennedy (2022) and Schuler and van der Laan (2025)), the EIF for a specific estimand can be used to construct an estimator that is doubly robust and attains the nonparametric efficiency bound (a nonparametric analog of the Cramer-Rao bound), at least asymptotically. Note that, as we will see below, the EIF does not always construct a doubly-robust estimator (Schuler and van der Laan, 2025), and the EIF may not exist for a particular estimand. In practice, one has to derive the EIF only for a novel estimand, and in many cases this can build closely on past work. Kennedy et al. (2017) provided a doubly-robust estimation method for causal exposure-response curves in observational data settings, which Hettinger et al. (2025a) then extended to a DiD framework, including conditions for “multiple robustness” which will be described in the following section. We further extend this multiply-robust estimator to a setting without control locations, which as described in Section 1 is another common type of quasi-experiment. This work contributes to the nascent application of EIF-based estimation techniques in quasi-experimental settings.
@@ 3.3 Estimation
The major difference between our estimator and that of Hettinger et al. (2025a) is that instead of modeling the pre-post difference in outcomes, we are just modeling the outcome $Y$ because we are using pre-intervention years rather than separate control locations. This does not substantively change the steps of the proofs in Hettinger et al. (2025a) to (1) derive the EIF and (2) show that the resulting estimator is multiply robust. A technical note is that the derivation of the EIF here refers to the integral of $ADT(\delta)$ to satisfy the pathwise differentiability requirement.
@@
$\xi(D,\mathbf{X},B,Y;\mu_{1},\pi_{D})=B*\left[\frac{f(D|B=1)}{\pi_{D}(D|\mathbf{X},B=1)}(Y-\mu_{1}(D,\mathbf{X}))+m(D|B=1)\right]$
$\tau(\mathbf{X},B,Y;\mu_{0},\pi_{B})=\frac{1}{P(B=1)}\left[\frac{(1-B)*\pi_{B}(\mathbf{X})}{(1-\pi_{B}(\mathbf{X}))}(Y-\mu_{0}(\mathbf{X}))+B*\mu_{0}(\mathbf{X})\right]$
@@
$f(D|B=1)=\int_{\mathbb{X}}\pi_{D}(D|B=1,\mathbf{x})dP(\mathbf{x}|B=1)$
@@
The estimation of the nuisance functions can be done with flexible modeling techniques subject to some constraints which prevent overfitting. For multiple robustness, $\mu_{1}$ and $\pi_{D}$ must either be fit using methods from uniformly bounded function classes with finite uniform entropy integrals (referred to as the Donsker class) or use cross-fitting / sample splitting, where the functions are estimated on one part of the data and used to predict in a held-out fold. The methods used to fit $\mu_{0}$ and $\pi_{B}$ just have to converge in probability to the truth (as $n\rightarrow\infty$).
@@ 3.4 Uncertainty Quantification
For uncertainty quantification, Hettinger et al. (2025a) derive a sandwich variance estimator, but mention various limitations of that approach, including that it relies on a normal approximation, that incorporating uncertainty in nuisance function estimation is challenging, and that extensions would be required to account for correlated data. To avoid these complications, they use a nonparametric bootstrap approach to obtain pointwise confidence intervals for the ADT. Rather than discrete bootstrapping, where observations are either included in a sample or left out entirely, they use weights drawn from an Exponential distribution to allow for small but nonzero inclusion, which helps to improve the observed support of the continuous exposure (in our notation, $D$) in the samples (Hettinger et al., 2025a). To account for temporal correlation across repeated observations, they sample one weight per location (as opposed to per individual observation). These weights are scaled to reflect the observed sample sizes in each treatment group.
@@
Note that our approach to addressing spatial correlation is distinct from the weighted block bootstrap approach in a working paper by Hettinger et al. (2025b), in which they sample a weight per block/neighborhood defined as all adjacent ZIP codes to a focal ZIP code, and then sum the weights assigned within each ZIP code from overlapping blocks. Described below, our approach has the advantages of (a) empirically estimating the residual spatial correlation and (b) avoiding artifacts/inconsistencies stemming from the use of location-areas with irregular sizes and/or shapes, such as two sites $m$ miles apart being considered spatially correlated in some places but not in other places, and ZIP codes with more adjacencies (shared borders) systematically receiving larger weights.
@@ 3.5 Case Study Data Application
In our case study analysis, we use the Ranger algorithm in the caret package (Kuhn and Max, 2008) in R to estimate the nuisance functions, using separate models for $\mu_{1},\mu_{0},\pi_{B},\pi_{D}$. Ranger is an efficient implementation of random forest, which has been shown to work well in tabular data settings like this (Grinsztajn et al., 2022), and also implicitly performs a type of variable selection (James et al., 2022), which is helpful for identifying which of the fixed effects, at the province or individual site level, is informative (Grinsztajn et al., 2022). However, because this algorithm does not conform to the finite uniform entropy integral condition used to prove multiple robustness, we use cross-fitting with k=10 sample splits for $\mu_{1}$ and $\pi_{D}$. (For $\mu_{0}$ and $\pi_{B}$ we do not do cross-fitting to reduce the computational intensity – as it is not needed for multiple robustness – instead carefully tuning hyperparameters as described below.) Note that Donsker class alternatives like Highly Adaptive Lasso (HAL) tend to be substantially more computationally intensive than Ranger, even with cross-fitting. Another important note is that cross-fitting in a spatiotemporal setting like this one is an active area of research. Splitting samples into folds by unit and by time period both have their disadvantages. Interestingly, in a variety of panel data simulations, Fuhr and Papies (2024) found violations of the cross-fitting independence assumption to be largely inconsequential for the accuracy of the effect estimates.
Note that while we referred to a common set of covariates ($\mathbf{X}$) throughout the previous sections, the set of covariates used in the two propensity score models is actually slightly smaller than that described in Section 3.1. The time (month) variable is not included in the model for $\pi_{B}$, because the set of covariates in a flexible machine learning model must not be able to perfectly predict the intervention (Heffernan et al., 2024) – in our case, time steps before and after 1/1/2018. Similarly, the dump site ID is not included in the model for $\pi_{D}$, because the port proximity value is unique to each site. The time variable is also not included in the model for $\pi_{D}$ because our port proximity values are static over time. Note that because of this, we sample observations by site into the cross-fitting folds (in addition to initial hyperparameter tuning).
@@
where $\Gamma$ is the gamma function, $K_{\nu}$ is the modified Bessel function of the second kind, and $\nu$ is an additional smoothness parameter which is optimized alongside $L$. The empirical spatial range was $\hat{L}=3.85$km, and the optimal smoothness parameter was $\hat{\nu}=5$. Figure 3 illustrates the resulting magnitude of correlation for sites at different distances apart.
Figure 3: The spatial correlation function (Matérn) fit to the residuals from the LLKR stage.
@@
For illustrative purposes, we compare the results from our multiply-robust EIF-based method to (i) a plain outcome regression, which is the same as that used inside the multiply-robust estimator, followed by regression on $D$ using LLKR, and (ii) a flexible but confounder-naive approach. The latter estimates $E[Y|D=\delta,B=1]-E[Y|B=0]$, where the first expectation uses LLKR to regress on $D$ and the second expectation is a sample mean. We also re-ran the analysis using 2017 and, separately, 2016 & 2017 as the “treated” period, to check for pre-trends / policy anticipation (whether the effect began before 2018). Conducting such tests is standard practice for quasi-experiments; Heffernan et al. (2024) demonstrated this technique while utilizing past years as controls.
This pre-trends validation approach connects to our stated assumptions in several ways. Clearly, anticipation of a policy being implemented would violate A1 (arrow of time). A5, on the other hand, would be violated if there was some unmeasured time-varying factor causing PM2.5 to be systematically different pre- and post-2018, in the absence of China’s ban. If the pre-trends analysis finds an effect of treatment in 2017 (prior to treatment initiation), this could also indicate that some important time-varying factors influencing PM2.5 are not being captured by our measured confounders, in violation of A5. (A policy implemented in 2017 could fit this criteria, as noted in our Discussion.) However, the absence of an effect in the pre-trends test only offers evidence of plausibility of A5 but does not formally validate it because we cannot test for unmeasured time-varying confounding between $B=0$ and $B=1$, instead relying on domain knowledge. Also, note that because our pre-trends test is estimating the ADT, it is inherently testing conditional ignorability in expectation.

### v4

@@ 1 Introduction
Our proposed method merges ideas from this literature on quasi-experiments with a “universal” intervention and ideas from the burgeoning literature on multiply-robust causal exposure-response curve estimation. Specifically, we build closely upon recent work by Hettinger et al. (2025b), who introduced a multiply-robust estimator for causal exposure-response curves within the difference-in-differences (DiD) framework. Derived from the efficient influence function or EIF (see Section 3.2), this estimator accounts for confounding both between treated and control units and across levels of exposure to the treatment / intervention, which can be thought of as a dosage. We propose a related EIF-based multiply-robust estimator for causal exposure-response curves in quasi-experimental designs with a universal intervention. We also propose an uncertainty quantification approach that accounts for both spatial and temporal correlation. We apply this method to remotely sensed data for the Indonesia plastic waste case study to estimate the change in air pollution at open dump sites due to increased plastic waste in the wake of China’s 2018 ban, as a function of a dump site’s port proximity.
@@ 2.1.1 Why Non-dump Sites Should Not Be Used as “Control” Locations
To motivate our methodological approach (in which we use past years as controls for future years at the GPW dump sites), briefly consider what trying to identify separate “control” locations would look like. Let us walk through several of the causal identifying assumptions made in Hettinger et al. (2025b), which are standard in DiD-type analyses.
@@
We extract these monthly PM2.5 values for 2012-2019 at the GPW sites, for each site taking the value of the grid cell in which the site centroid lies. The average PM2.5 concentration at the dump sites in 2012-2017 (prior to China’s ban) was 25.7 $\mu g/m^{3}$.
@@ 3.1 Setup and Assumptions
Let $\mathcal{O}$ be the observed data, containing: $Y$ the outcome (PM2.5), $D$ the dose of exposure (port proximity), and $\mathbf{X}$ the covariates. The latter includes the population living within 1km of the site, monthly averaged meteorology (temperature, humidity, precipitation, pressure, and wind speed), and the number of active fires detected in the province (that month) which did not overlap with any GPW site. To help capture spatial heterogeneity, we also include fixed effects for site ID and province in our model; the spatial boundaries for the latter are shown in Figure S3. This is considered a reasonable approach in settings with many units where only few of these are important/informative (Fuhr and Papies, 2024) and when both bigger-scale (e.g., province level) and smaller-scale (individual dump site level) heterogeneity are likely at play. Both of these characteristics are present in our case study, as shown by a complementary exploratory analysis (described in Section 4.2) illustrating that the distribution of the number of large fires (detectable by VIIRS) occurring at each dump site is highly skewed, with some clustering by province. Lastly, to account for longer-term trends, we include time t=1,,96t=1,...,96 to represent the months between 2012 and 2019, inclusive. Note that the seasonality illustrated in Figure 1 is encoded in the meteorological covariates and the number of active fires detected in the province which did not overlap with any GPW site, both of which are time-varying.
@@
In order to avoid structural ill-definition, we also introduce an indicator for $t\geq T_{0}$, where $T_{0}$ is the treatment initiation time. In practice, $T_{0}$ is 1/1/2018 for all sites, but conceptually this could have occurred at any point in our study period, such that any time point could have been untreated. When $T_{0}$ is considered random, A5 (below) is a direct analog of the conventional conditional ignorability assumption used in cross-sectional data contexts, adapted to the time series setting.
Our aim is to estimate a causal exposure-response curve quantifying the causal effect on Indonesian air quality of China’s 2018 plastic waste import ban at different levels of port proximity, $\delta$. In the DiD setting, Hettinger et al. (2025b) terms this the “Average Dose Effect on the Treated”, or ADT. While noting again that our formulation is different, we still see fit to call our effect of interest the ADT:
@@
An important note about the ADT is how it differs from a more commonly-used Conditional Average Treatment Effect, $CATE(\delta):=E[Y_{it}(B=1)-Y_{it}(B=0)|D_{it}=\delta]$. First, recall that $D$ is undefined when $B=0$, not yet having been ‘activated’. Second, foreshadowing the following sections, the ADT approach enables interpreting the final effect curve in terms of port proximity versus in terms of port proximity-possibly-confounded-by-other-factors. It does so by targeting confounding with respect to both $D$ and $B$, estimating a generalized propensity score for the former and a binary propensity score for the latter, in addition to modeling the outcome. A traditional doubly-robust estimator for the CATE would only use the binary propensity score.
@@
If $\mathbf{X}^{B}$ is the covariates excluding the time indicator $t$ and $\mathbf{X}^{D}$ is the covariates excluding both the dump site ID and the time indicator $t$, then we require
@@
$\>\>\>\>\text{(i) }\epsilon\leq\pi_{B}(\mathbf{x}^{B})<1\>\forall\>\mathbf{x}^{B}\in\mathbb{X}^{B},\>\>\>\>\>\>\>\>\pi_{B}(\mathbf{x}^{B})=P(t\geq T_{0}|\mathbf{x}^{B})$
$\>\>\>\>\text{(ii) }\epsilon\leq\pi_{D}(\delta|\mathbf{x}^{D},B=1)\>\forall\>\mathbf{x}^{D}\in\mathbb{X}^{D}|B=1,\delta\in\mathbb{D},\>\>\>\>\pi_{D}(\delta|\mathbf{x^{D}},B=1)=p(D=\delta|B=1,\mathbf{X}^{D}=\mathbf{x}^{D})$
While $P(t\geq T_{0}|\mathbf{x}^{B})$ does not represent a feasible data generating process, this formulation highlights that there must not be separation in the covariate space between the pre- and post-intervention time periods. Flexibly estimating $\pi_{B}$ characterizes the extent of this overlap, rather than attempting to capture a true data generating process.
@@
This means that there is no unmeasured time-varying confounding – no unmeasured factor causing air pollution to be systematically different in the periods before and after China’s ban, after accounting for confounding factors. Assembling an adequate set of covariates to justify this assumption relies on domain knowledge. Its plausibility can be strengthened by adjusting for seasonal and/or long-term trends, e.g., through inclusion of a time variable in the adjustment set. Note that while China’s ban being enacted should not affect confounders like meteorology and population density in Indonesia, we must pay careful attention to any other policies with the potential to make air pollution, conditional on covariates, systematically different in the $B=0$ and $B=1$ time periods – see discussion of this in Section 5.
@@
Conditional ignorability (in both A5 and A6) is a stronger assumption than parallel trends. Any method for treatment effect evaluation in the absence of a control group must make assumptions to rule out unforecastable shocks that affect all outcomes in the post-treatment period (Botosaru et al., 2026). Similarly, for a continuous treatment, while parallel trends would adjust for baseline differences between units receiving different doses, we must assume that our covariate set is able to explain these baseline differences. For our application, the alternative of assuming that some locations represent the control condition post-treatment (China’s ban) would likely be inaccurate, as described in Section 2.1.1.
Note that while we believe it is reasonable to assume conditional ignorability on distributions here given our domain knowledge, all that is needed for identifiability of our estimand is (a weaker) conditional ignorability in expectations, as is shown below.
@@ 3.2 Motivations for Efficient Influence Function-Based Estimation
where $f_{\phi}(D)$ is a flexible (nonparametric) function, with parameters $\phi$. Note that a simplified version of this, where $f_{\phi}(D)=\phi D$, has been extensively analyzed – for instance see Callaway et al. (2024). This approach has several drawbacks. To begin, even if we allowed a flexible relationship between $D$ and $Y$, we would not be allowing for flexible relationships between $\mathbf{X}$ and $Y$, nor interactions between the components of $\mathbf{X}$. Importantly, we would not be addressing confounding between the dose $D$ and trends in the outcome – for instance, if the overall (or seasonal) trend in air pollution is affected by population density, which is static in our dataset and which we know to be correlated with $D$ (see Section 4.2). To address this, we could make a much more complicated linear regression (e.g., by using splines and more interaction terms), or use a nonparametric machine learning algorithm to model $Y$ given $D,\mathbf{X},B$. However, this outcome model might still be misspecified. This motivates the search for a doubly-robust estimator, which incorporates both a model for the outcome and a model for the propensity score (the probability of receiving treatment), which explicitly addresses confounding, and will produce results that are consistent (reliable in a probabilistic sense) if at least one of these models is correctly specified.
We turn to estimators based on the the Efficient Influence Function (EIF). Without going into the technical details (interested readers should see Kennedy (2022) and Schuler and van der Laan (2025)), the EIF for a specific estimand can be used to construct an estimator that is doubly robust and attains the nonparametric efficiency bound (a nonparametric analog of the Cramer-Rao bound), at least asymptotically. Note that, as we will see below, the EIF does not always construct a doubly-robust estimator (Schuler and van der Laan, 2025), and the EIF may not exist for a particular estimand. In practice, one has to derive the EIF only for a novel estimand, and in many cases this can build closely on past work. Kennedy et al. (2017) provided a doubly-robust estimation method for causal exposure-response curves in observational data settings, which Hettinger et al. (2025b) then extended to a DiD framework, including conditions for “multiple robustness” which will be described in the following section. We further extend this multiply-robust estimator to a setting without control locations, which as described in Section 1 is another common type of quasi-experiment. This work contributes to the nascent application of EIF-based estimation techniques in quasi-experimental settings.
@@ 3.3 Estimation
The major difference between our estimator and that of Hettinger et al. (2025b) is that instead of modeling the pre-post difference in outcomes, we are just modeling the outcome $Y$ because we are using pre-intervention years rather than separate control locations. This does not substantively change the steps of the proofs in Hettinger et al. (2025b) to (1) derive the EIF and (2) show that the resulting estimator is multiply robust. A technical note is that the derivation of the EIF here refers to the integral of $ADT(\delta)$ to satisfy the pathwise differentiability requirement.
@@
$\xi(D,\mathbf{X},B,Y;\mu_{1},\pi_{D})=B*\left[\frac{f(D|B=1)}{\pi_{D}(D|\mathbf{X}^{D},B=1)}(Y-\mu_{1}(D,\mathbf{X}))+m(D|B=1)\right]$
$\tau(\mathbf{X},B,Y;\mu_{0},\pi_{B})=\frac{1}{P(B=1)}\left[\frac{(1-B)*\pi_{B}(\mathbf{X}^{B})}{(1-\pi_{B}(\mathbf{X}^{B}))}(Y-\mu_{0}(\mathbf{X}))+B*\mu_{0}(\mathbf{X})\right]$
@@
$f(D|B=1)=\int_{\mathbb{X}}\pi_{D}(D|B=1,\mathbf{x}^{D})dP(\mathbf{x}^{D}|B=1)$
@@
The estimation of the nuisance functions can be done with flexible modeling techniques subject to some constraints which prevent overfitting. For multiple robustness, $\mu_{1}$ and $\pi_{D}$ must either be fit using methods from uniformly bounded function classes with finite uniform entropy integrals (referred to as the Donsker class) or use cross-fitting / sample splitting, where the functions are estimated on one part of the data and used to predict in a held-out fold. The methods used to fit $\mu_{0}$ and $\pi_{B}$ just have to converge in probability to the truth (as $n\rightarrow\infty$). Recent work on estimating treatment effects in the absence of a control group (Botosaru et al., 2026) emphasizes the importance of unbiasedness rather than predictive accuracy when forecasting the counterfactual for the treated based on pre-treatment observations. (Other aspects of Botosaru et al. (2026) do not apply in our setting because we have a much longer time series and complex covariate space that cannot easily be forecasted with basis functions.)
@@ 3.4 Uncertainty Quantification
For uncertainty quantification, Hettinger et al. (2025b) derive a sandwich variance estimator, but mention various limitations of that approach, including that it relies on a normal approximation, that incorporating uncertainty in nuisance function estimation is challenging, and that extensions would be required to account for correlated data. To avoid these complications, they use a nonparametric bootstrap approach to obtain pointwise confidence intervals for the ADT. Rather than discrete bootstrapping, where observations are either included in a sample or left out entirely, they use weights drawn from an Exponential distribution to allow for small but nonzero inclusion, which helps to improve the observed support of the continuous exposure (in our notation, $D$) in the samples (Hettinger et al., 2025b). To account for temporal correlation across repeated observations, they sample one weight per location (as opposed to per individual observation). These weights are scaled to reflect the observed sample sizes in each treatment group.
@@
Note that our approach to addressing spatial correlation is distinct from the weighted block bootstrap approach in a working paper by Hettinger et al. (2025a), in which they sample a weight per block/neighborhood defined as all adjacent ZIP codes to a focal ZIP code, and then sum the weights assigned within each ZIP code from overlapping blocks. Described below, our approach has the advantages of (a) empirically estimating the residual spatial correlation and (b) avoiding artifacts/inconsistencies stemming from the use of location-areas with irregular sizes and/or shapes, such as two sites $m$ miles apart being considered spatially correlated in some places but not in other places, and ZIP codes with more adjacencies (shared borders) systematically receiving larger weights.
@@ 3.5 Case Study Data Application
In our case study analysis, we use the Ranger algorithm in the caret package (Kuhn and Max, 2008) in R to estimate the nuisance functions, using separate models for $\mu_{1},\mu_{0},\pi_{B},\pi_{D}$. Ranger is an efficient implementation of random forest, which has been shown to work well in tabular data settings like this (Grinsztajn et al., 2022), and also implicitly performs a type of variable selection (James et al., 2022), which is helpful for identifying which of the fixed effects, at the province or individual site level, is informative (Grinsztajn et al., 2022). However, because this algorithm does not conform to the finite uniform entropy integral condition used to prove multiple robustness, we use cross-fitting with k=10 sample splits for $\mu_{1}$ and $\pi_{D}$. (For $\mu_{0}$ and $\pi_{B}$ we do not do cross-fitting to reduce the computational intensity – as it is not needed for multiple robustness – instead carefully tuning hyperparameters as described below. For applications with simpler covariate spaces and/or shorter pre-treatment time series, a parametric or semi-parametric model to forecast the counterfactual may be preferred (Botosaru et al., 2026).) Note that Donsker class alternatives like Highly Adaptive Lasso (HAL) tend to be substantially more computationally intensive than Ranger, even with cross-fitting. Another important note is that cross-fitting in a spatiotemporal setting like this one is an active area of research. Splitting samples into folds by unit and by time period both have their disadvantages. Interestingly, in a variety of panel data simulations, Fuhr and Papies (2024) found violations of the cross-fitting independence assumption to be largely inconsequential for the accuracy of the effect estimates.
Recall from Section 3.1 that the set of covariates used in the two propensity score models excludes the time (month) indicator $t$. For $\pi_{B}$, this is because the set of covariates in a flexible machine learning model must not be able to perfectly predict the intervention (Heffernan et al., 2024) – in our case, time steps before and after 1/1/2018. For $\pi_{D}$, this is because our port proximity values are static within the $B=1$ time period. Additionally, the covariates used to estimate $\pi_{D}$ do not include the dump site ID because the port proximity value is unique to each site. Because of this, we sample observations by site into the cross-fitting folds (in addition to initial hyperparameter tuning).
@@
where $\Gamma$ is the gamma function, $K_{\nu}$ is the modified Bessel function of the second kind, and $\nu$ is an additional smoothness parameter which is optimized alongside $L$.
The empirical spatial range was $\hat{L}=3.85$km, and the optimal smoothness parameter was $\hat{\nu}=5$. Figure 3 illustrates the resulting magnitude of correlation for sites at different distances apart.
Figure 3: The spatial correlation function (Matérn) fit to the residuals from the LLKR stage; $\hat{L}=3.85$km and $\hat{\nu}=5$.
@@
For illustrative purposes, we compare the results from our multiply-robust EIF-based method to (i) a plain outcome regression, which is the same as that used inside the multiply-robust estimator, followed by regression on $D$ using LLKR, and (ii) a flexible but confounder-naive approach. The latter estimates $E[Y|D=\delta,B=1]-E[Y|B=0]$, where the first expectation uses LLKR to regress on $D$ and the second expectation is a sample mean.
We also re-ran the analysis using 2017 and, separately, 2016 & 2017 as the “treated” period, to check for pre-trends / policy anticipation (whether the effect began before 2018). Conducting such tests is standard practice for quasi-experiments (Bilinski and Hatfield, 2026; Roth, 2022); Heffernan et al. (2024) demonstrated this technique while utilizing past years as controls, in other words, utilizing a placebo in time. Note that this kind of pre-trends / placebo testing is technically assessing the null hypothesis of no pre-trends; for more formal evaluation, it is preferred to test whether the magnitude of the pre-trend effect is sufficiently small as to not invalidate the estimated treatment effect (Bilinski and Hatfield, 2026). Future work might explore development of such tests for settings like ours.
This pre-trends validation approach connects to our stated assumptions in several ways. Clearly, anticipation of a policy being implemented would violate A1 (arrow of time). If the pre-trends analysis finds an effect of treatment in 2017 (prior to treatment initiation), this could also indicate that some important time-varying factors influencing PM2.5 are not being captured by our measured confounders, in violation of A5. (A policy implemented in 2017 could fit this criteria, as noted in our Discussion.) However, the absence of an effect in the pre-trends test only offers evidence of plausibility of A5 but does not formally validate it because we cannot test for unmeasured time-varying confounding between $B=0$ and $B=1$, instead relying on domain knowledge. Also, note that because our pre-trends test is estimating the ADT, it is inherently testing conditional ignorability in expectation.